OpenStudy (anonymous):

y<_ x + 4 2x+ y <_ -4 What is the graph of the system Connexus 10th grade lesson 7 unit 6 help

OpenStudy (ybarrap):

Is the the problem: y <= x + 4 2x + y <= -4 In latex (you may not be able to see this): $$ y \le x + 4\\ 2x + y \le -4 $$